OpenStudy (anonymous):

pls help how to do this one maths t=12n

OpenStudy (curry):

t/12=n

OpenStudy (anonymous):

how do u get this pls explain

OpenStudy (curry):

well think about it 12 times n equals t therefore t/12 = n and or u can try plugging in numbers and when u want to isolate you always do the reverse of operation when ur doing something to the other side

OpenStudy (curry):

please leave medal if it helped

OpenStudy (anonymous):

i m new to this one how do i give u medal

OpenStudy (curry):

click on good answer

OpenStudy (curry):

and then hover over my picture and click become fan

OpenStudy (anonymous):

hi can u show me how to do this t=6-3square n